sin -1 x + sin -1 2x = ∏/3

sin-1x + sin-12x = ∏/3

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Aby

  -1≤ 2x≤1

  -1/2 ≤x≤1/2

now

sin-1x + sin-12x = ∏/3

 sin-12x = ∏/3  -sin-1x

take sin

   2x = sin∏/3 cos sin-1x  -cos∏/3  sinsin-1x

        =√3 /2 √(1-x2)  -1/2 x

 or  5x/2  =√3 /2 √(1-x2)

square

  25x2 =3 (1-x2)

  x2 =3/28

 x= =±√3/28

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin